Find the general solution of the differential equation y" + 16y = 2 sec(4t) tan(4t)

Select one:

O none of the given answers is true

y = ci cos(4t) + c2 sin(4t) + ,t cos(4t) - sin(4t) In | cos(4t) |


y = ci cos(4t) + c2 sin(4t) - t cos(4t) + sin(4t) In | cos(4t) |

y = ci cos(4t) + c2 sin (4t) -

7t sin(4t) - - cos(4t) In | sin(4t) |

y = ci cos(4t) + C2 sin(4t) + ,t sin(4t) + - cos(4t) In | sin(4t)|



1
Expert's answer
2021-08-17T10:36:50-0400

Corresponding homogeneous differential equation


"y''+16y=0"

Characteristic equation


"r^2+16=0"

"r=\pm4i"

The general solution of the homogeneous differential equation is


"y=C_1\cos(4t)+C_2\sin(4t)"

Use the variation of parameters


"W(y_1, y_2)=\begin{vmatrix}
 y_1 & y_2 \\
 y_1' & y_2'
\end{vmatrix}=\begin{vmatrix}
 \cos(4t) & \sin(4t) \\
 -4\sin(4t) & 4\cos(4t)
\end{vmatrix}=4"

Find the partial solution


"y_p=-\cos(4t)\int\dfrac{\sin(4t)\sec(4t)\tan(4t)}{4}dt"

"+\sin(4x)\int\dfrac{\cos(4x)\sec(x)\tan(x)}{4}dx"

"\int\dfrac{\sin(4t)\sec(4t)\tan(4t)}{4}dt=\int\dfrac{\tan^2(4t)}{4}dt"

"=\int\dfrac{\sec^2(4t)-1}{4}dt=\dfrac{\tan(4t)}{16}-\dfrac{t}{4}-C_3"



"\int\dfrac{\cos(4t)\sec(4t)\tan(4t)}{4}dt=\int\dfrac{\tan(4t)}{4}dt"

"=\int\dfrac{\sin(4t)}{4\cos(4t)}dt=-\dfrac{\ln|\cos(4t)|}{16}+C_4"

"y_p=-\cos(4t)\dfrac{\tan(4t)}{16}+\dfrac{t}{4}\cos(4t)-C_3\cos(4t)"

"-\sin(4t)\dfrac{\ln|\cos(4t)|}{16}+C_4\sin(4t)"

The general solution of the given differential equation is


"y=c_1\cos(4t)+c_2\sin(4t)"

"-\dfrac{\sin(4t)}{16}+\dfrac{t}{4}\cos(4t)-\sin(4t)\dfrac{\ln|\cos(4t)|}{16}"